# Multilayered Intelligent Reflecting Surface for Long-Range Underwater Acoustic Communication

### v1

@@ IV Hardware Architecture
This section outlines the hardware design of the UA-RIS. The multilayer structure of the acoustic reflector is introduced first. Subsequently, we present the circuit designed to generate in-phase and quadrature-reflected waves for effective IQ modulation.
@@ IV-A Multilayered Acoustic Reflector
In ML-ARIS, we leverage the acoustic coupling property inherent in mechanical waves to implement IQ modulation within a single reflector. This property facilitates the transfer of mechanical vibrations across the medium, enabling energy propagation into subsequent layers. Unlike electromagnetic fields, where a conventional RF reflector acts as a conductive boundary that prevents wave penetration beyond the initial layer, the multilayer structure proposed in this work is specific to UA-RIS and fundamentally differs from existing solutions developed for EM-based applications.
@@
The structure of the acoustic reflector designed for the ML-ARIS system is illustrated in Fig. 2. This reflector employs a Tonpilz-style configuration to achieve efficient electroacoustic conversion in the mid-frequency band. Within the reflector, a head mass composed of aluminum alloy promotes effective oscillation along the central axis while minimizing stress on other components. A stack of PZT ceramic disks is sandwiched between the head and back masses, serving as the core elements that control the reflected waves. To enhance acoustic coupling efficiency, the PZT stack is held under a constant compressive preload by a prestressing rod. At the rear, a tail mass reflects backward-traveling vibrations, ensuring that the reflected waves propagate forward.
@@ IV-B Control Circuit Design
Fig. 3 illustrates the control circuit developed for a two-layer acoustic reflector. The system employs two identical load networks, each composed of four primary subnetworks: a matching network, a resistive network, a capacitive network, and an inductive network. The TCA9555 I/O extender, interfacing with the ATmega256RFR2 MCU via the I2C protocol, governs the configuration of each load network. This arrangement enables precise manipulation of both the phase and amplitude of the waves reflected by the corresponding PZT disk, thereby facilitating effective IQ modulation.
@@
The matching network aligns the PZT disk impedance to a specified real value $Z_{0}$, adapting environmental variations as will be detailed in Section V. The resistive network, which includes an AD8403AR50 programmable potentiometer, provides the in-phase component required for IQ modulation. When the pair of n-type MOS (NMOS) transistors in the resistive network is activated by the I/O extender, the potentiometer is engaged with the matching circuit. Under these conditions, the reflection coefficient of PZT disk 1 is given by
@@
As indicated by (2), when the load is a resistive component, $\Gamma$ is a real number, then the phases of the reflected and incident waves will be identical or reverse. By adjusting $R_{L}$ through programming the potentiometer, we can dynamically manipulate the amplitude of the in-phase component.
The capacitive and inductive networks are designed to generate the quadrature component for IQ modulation. Each subnetwork comprises three stages. Connecting different stages of the subnetwork to the matching circuit yields a reflection coefficient with a consistent phase of --90° (capacitive load) or 90° (inductive load) and three selectable amplitudes increasing stepwise from 0.3 to 0.9.
IV-C How IQ Modulation Works with ML-ARIS
[…]
@@ IV-C How IQ Modulation Works with ML-ARIS
As indicated in (3), the first term is the in-phase component, maintaining the same phase as the incident wave, while the second term is the quadrature component, which introduces a --90° phase shift relative to the incident wave.
@@
It should be noted that in a conventional Tonpilz-style acoustic transmitter, multiple PZT disks are connected in parallel and driven by the same electrical signal, resulting in uniform mechanical deformation across all disks. In contrast, for the multilayered reflector proposed here, the mechanical deformation varies among layers because the phases of the electric fields reflected by the load networks of adjacent PZT disks are orthogonal. Consequently, in practical implementations, the reflection signals generated within different layers exhibit a certain degree of coupling rather than being entirely independent. The influence of this coupling on the accuracy of the reflected waves will be examined through simulations and tank experiments, as detailed in Section VI.
@@ V Enhanced Matching Network
This section introduces an enhanced matching network designed to improve the adaptability of ML-ARIS across various environments. Initially, we analyze the necessity for a new matching network, followed by a detailed explanation of a simplified circuit model for underwater acoustic reflectors. Subsequently, we present the architecture of the new matching network and explore the methods employed to optimize its performance under diverse conditions.
V-A Necessity for a New Matching Network
In RF-IRS systems, the impedance of the reflector changes exclusively with frequency. Under these conditions, a wide-band matching circuit proves highly effective in aligning the impedance of the reflector with that of the load. However, for underwater acoustic reflectors constructed from PZT material, the impedance changes not only with frequency but also in response to temperature of the water, incident angle, and deployment depth. To ensure efficient operation of the ML-ARIS across various application scenarios, the matching circuit must be wide-band and adaptable to environmental changes.
V-A1 Water temperature & depth
In Fig. 4, an Agilent E5071C vector network analyzer (VNA) is employed to investigate how the impedance of a PZT disk varies with signal frequency and water temperature. The structure of the reflector used in the test has been introduced in Fig. 2. The resonate frequencies of all disks are around 28.2 kHz at 20°C. In the figure, the symbols $R$ and $X$ denote the real and imaginary components of the impedance at the first PZT disk, respectively.
@@ V-A1 Water temperature & depth
As illustrated in Fig. 4, both $R$ and $X$ exhibit substantial changes at the resonant frequency with varying temperatures. For instance, at a water temperature of 9°C and a signal frequency of 28.2 kHz, $R$ is 678 $\Omega$ and $X$ is 142 $\Omega$. As the temperature rises to 22°C, $R$ decreases to 494 $\Omega$, and $X$ shifts to --247 $\Omega$. As a result, with the water temperature rising from 9°C to 22°C, the magnitude of the PZT disk impedance at 28.2 kHz increases by 20%, and the reactance of the impedance transitions from capacitive to inductive.
In addition to the variation of temperature, the impedance of the PZT material also changes with water depth. Specifically, the capacitance of PZT, which depends on their geometry and dielectric properties, is particularly sensitive to these changes. Increased water depth results in heightened pressure, which imposes mechanical stress on the material. This stress compresses the thickness of the material, consequently increasing its capacitance, which in turn leads to a reduction in the imaginary component of the impedance.
V-A2 Incident angle
For a typical RF-RIS operating in free space, the direction of incoming waves does not significantly alter the electrical reflector’s input impedance. In contrast, for an acoustic reflector in an underwater environment, the impedance varies noticeably with the incident angle due to changes in boundary conditions and how sound couples into or out of the piezoelectric material at different angles.
Moreover, the variation in impedance with incident angle may differ across the individual PZT layers. This phenomenon can be validated through a COMSOL multiphysics simulator. Specifically, in the simulation, a reflector comprising four PZT layers is constructed. Each layer measures 6 cm×$\times$0.48 cm and is separated by an acrylic plastic layer 0.2 mm in thickness. The incident signal is a monochromatic plane wave with a frequency of 41 kHz and an intensity of 1 Pa.
@@ V-A2 Incident angle
In Fig. 5, we vary the direction of incoming waves and measure the resistance and reactance for the first two layers of the PZT stack. The impedance in this figure is calculated from the ratio of the open-circuit voltage to the short-circuit current induced by the incoming waves. As illustrated, both the resistance and reactance of the PZT impedance demonstrate non-monotonic variations with changes in the incident angle. For example, at an incident angle of 90°, the resistance and reactance for PZT layer 1 are 8.9 $\Omega$ and --43.5 $\Omega$, respectively. When the incident angle is reduced to 45°, the resistance increases to 11.5 $\Omega$, a 29.2% increase from the initial value. Reducing the angle further to 30°, the reactance decreases to --46.1 $\Omega$, a 6% reduction from the 90° incident angle.
For PZT layer 2, at an incident angle of 90°, the resistance and reactance are 9.9 $\Omega$ and --42.7 $\Omega$, respectively. Changing the angle to 45° leads to an increase in resistance to 12.7 $\Omega$, marking a 28.3% rise from the 90°situation. When the angle is further decreased to 30°, the reactance rises to --38.1 $\Omega$, showing a 9.7% increase compared to the 90°scenario.
@@
Figure 6: Stress variation across 4 layers of the PZT stack, measured at 380 $\mu$s. (a) Incident angle: 90°. (b) Incident angle: 45°.
In Fig. 6, we present the stress distribution across four layers of the PZT stack when stimulated by monochromatic plane waves incident from two different directions at 380 $\mu$s. The mechanical wave propagation speed in PZT material is approximately 3900 m/s, and the total thickness of the four-layer stack corresponds to about 0.2 wavelengths. Under these conditions, it can be reasonably assumed that the mechanical wavefront reaches all layers simultaneously. The red dashed line in the figure indicates the maximum instantaneous stress within the stack. As shown in the figure, when the incident angle is 90°, the stress distribution across different layers is similar, and the red dashed line appears nearly straight. However, when the incident angle decreases to 45°, the stress distribution differs significantly among the layers, causing the red dashed line to form a pronounced arc.
@@ V-C Impedance Estimation
In $\mathbf{Z_{R}}$, let $Z^{max}_{R}(f_{r})$ be the element with the maximum magnitude at the resonate frequency $f_{r}$. The parameters used in (7) to fit $Z^{max}_{R}(f)$ are represented by $R^{\alpha}_{E}$, $C^{\alpha}_{E}$, and $\left[\text{Re}^{\prime}({Z_{S}})\right]^{\alpha}$. In contrast, the impedance with the maximum magnitude at $f_{r}$ is represented by $Z^{min}_{R}(f_{r})$; the parameters used to fit $Z^{min}_{R}(f)$ are denoted as $R^{\beta}_{E}$, $C^{\beta}_{E}$, and $\left[\text{Re}^{\prime}({Z_{S}})\right]^{\beta}$, where $R^{\alpha}_{E}\!<\!R^{\beta}_{E}$, $C^{\alpha}_{E}\!<\!C^{\beta}_{E}$, and $\left[\text{Re}^{\prime}({Z_{S}})\right]^{\alpha}\!>\!\left[\text{Re}^{\prime}({Z_{S}})\right]^{\beta}$.
@@
$\left\{\begin{array}[]{lll}\vspace{0.1cm}\!\!\!\!R^{i}_{E}=R^{\beta}_{E}-\displaystyle\frac{(i-1)\!\times\!\!\left(R^{\beta}_{E}-R^{\alpha}_{E}\right)}{N_{d}-1},\\ \vspace{0.1cm}\!\!\!\!C^{i}_{E}=C^{\beta}_{E}-\displaystyle\frac{(i-1)\!\times\!\!\left(C^{\beta}_{E}-C^{\alpha}_{E}\right)}{N_{d}-1},\\ \!\!\!\!\left[\text{Re}^{\prime}\!({Z_{S}})\right]^{\!i}\!\!=\!\left[\text{Re}^{\prime}\!({Z_{S}})\right]^{\!\beta}\!\!+\!\displaystyle\frac{\!(i\!-\!1)\!\!\times\!\!\left(\!\left[\text{Re}^{\prime}\!({Z_{S}})\right]^{\alpha}\!\!\!-\!\!\left[\text{Re}^{\prime}\!({Z_{S}})\right]^{\beta}\right)}{N_{d}-1},\\ \end{array}\right.$ (8)
@@ V-D1 Architecture
In Fig. 8, we illustrate the architecture of the enhanced matching circuit developed for the first layer of the PZT stack in one of the acoustic reflectors. This circuit features a three-tier subnetwork structure, each tier consisting of a high-pass L-type matching circuit. The interconnection between tiers is managed by NMOS transistors $\text{SW}_{1}$ and $\text{SW}_{2}$, which are controlled by the MCU. It is crucial to note that, the operation of the tiers is not independent but cascaded; for instance, activating tier $i$ requires that tiers 1 through $i\!-\!1$ also be activated to achieve the optimal performance.
@@ V-D2 Optimization
$\begin{array}[]{lll}\textbf{\emph{P1}}\;\;\vspace{0.1cm}\underset{C_{M_{1}},\,L_{M_{1}}}{\mathrm{arg\,min}}\,\displaystyle{\sum^{3}_{i=1}}\,\displaystyle{\sum^{f_{H}}_{f=f_{L}}}\,\absolutevalue{\displaystyle\frac{Z_{0}-Z^{i}_{M_{1}}(f)}{Z_{0}+Z^{i}_{M_{1}}(f)}}^{3},\\ \begin{array}[]{lll}\vspace{0.05cm}\textbf{s.t.}&\textbf{\emph{C1:}}\quad C_{M_{1}}>0,\\ &\textbf{\emph{C2:}}\quad L_{M_{1}}>0.\end{array}\end{array}$ (10)
@@ VI Performance Evaluation
This section validates the feasibility of ML-ARIS through simulation and experimentation. Initially, we test the accuracy of the circuit model developed for the PZT reflector. Subsequently, the performance of the enhanced matching circuit is assessed. We then present experimental results from both tank and lake environments, which confirm the operational viability of the system in real-world applications.
VI-A Verification of IQ Modulation with Multilayered Structure
VI-A1 Experimental settings
In Fig. 9, tank experiments are performed to validate the feasibility of IQ modulation using ML-ARIS. In the tests, two BTech Acoustic BT-2RCL [25] omnidirectional acoustic transducers serve as the transmitter and receiver, respectively. The transmitter is driven by a Siglent 1062X arbitrary waveform generator emitting a burst sinusoidal signal consisting of 2,500 cycles. The frequency and peak-to-peak voltage of the source signal are 43.8 kHz and 20 V, respectively. The receiver is connected to a PicoScope 4224A oscilloscope for data acquisition.

### v2

@@ IV Hardware Architecture
This section outlines the hardware design of the UA-RIS. The multilayer structure of the acoustic reflector is introduced first. Subsequently, we present the circuit designed to generate in-phase and quadrature-reflected waves for effective synthetic reflection.
@@ IV-A Multilayered Acoustic Reflector
In ML-ARIS, we leverage the acoustic coupling property inherent in mechanical waves to implement synthetic reflection within a single reflector. This property facilitates the transfer of mechanical vibrations across the medium, enabling energy propagation into subsequent layers. Unlike electromagnetic fields, where a conventional RF reflector acts as a conductive boundary that prevents wave penetration beyond the initial layer, the multilayer structure proposed in this work is specific to UA-RIS and fundamentally differs from existing solutions developed for EM-based applications.
@@
The structure of the acoustic reflector designed for the ML-ARIS system is illustrated in Fig. 3. This reflector employs a Tonpilz-style configuration to achieve efficient electroacoustic conversion in the mid-frequency band [28, 6]. Within the reflector, a head mass composed of aluminum alloy promotes effective oscillation along the central axis while minimizing stress on other components. A stack of PZT ceramic disks is sandwiched between the head and back masses, serving as the core elements that control the reflected waves. To enhance acoustic coupling efficiency, the PZT stack is held under a constant compressive preload by a prestressing rod. At the rear, a tail mass reflects backward-traveling vibrations, ensuring that the reflected waves propagate forward.
@@ IV-B Control Circuit Design
Fig. 4 illustrates the control circuit developed for a two-layer acoustic reflector. The system employs two identical load networks, each composed of four primary subnetworks: a matching network, a resistive network, a capacitive network, and an inductive network. The TCA9555 I/O extender, interfacing with the ATmega256RFR2 MCU via the I2C protocol, governs the configuration of each load network. This arrangement enables precise manipulation of both the phase and amplitude of the waves reflected by the corresponding PZT disk, thereby facilitating effective synthetic reflection.
@@
The matching network aligns the PZT disk impedance to a specified real value $Z_{0}$, adapting environmental variations as will be detailed in Section V. The resistive network, which includes an AD8403AR50 programmable potentiometer, provides the in-phase component required for synthetic reflection. When the pair of n-type MOS (NMOS) transistors in the resistive network is activated by the I/O extender, the potentiometer is engaged with the matching circuit. Under these conditions, the reflection coefficient of PZT disk 1 is given by
@@
As indicated by (2), when the load is a resistive component, $\Gamma$ is a real number; therefore, the phases of the reflected and incident waves will be identical or reverse. By adjusting $R_{L}$ through programming the potentiometer, we can dynamically manipulate the amplitude of the in-phase component.
The capacitive and inductive networks are designed to generate the quadrature component for synthetic reflection. Each subnetwork comprises three stages. Connecting different stages of the subnetwork to the matching circuit yields a reflection coefficient with a consistent phase of -90 (capacitive load) or 90 (inductive load) and three selectable amplitudes increasing stepwise from 0.3 to 0.9.
IV-C How Synthetic Reflection Works with ML-ARIS
Consider an incident wave represented by $\cos(\omega t\!+\!\phi_{in})$, where $\omega$ is the angular frequency, $t$ is time, and $\phi_{in}$ is the initial phase of the incident wave. The desired reflected wave is given by $A_{r}\cos(\omega t\!+\!\phi_{in}\!+\!\phi_{r})$, with $A_{r}\!\in\![\,0,1\,]$ and $\phi_{r}\!\in\![-\pi,\pi]$ denoting the amplitude and phase shift of the reflected wave, respectively. Using trigonometric identities, the reflected wave can be expressed as
@@ IV-C How Synthetic Reflection Works with ML-ARIS
As indicated in (3), the first term is the in-phase component, maintaining the same phase as the incident wave, while the second term is the quadrature component, which introduces a -90 phase shift relative to the incident wave.
@@
In a conventional Tonpilz-style acoustic transmitter, multiple PZT disks are connected in parallel and driven by the same electrical signal, resulting in uniform mechanical deformation across all disks. In contrast, for the multilayered reflector proposed here, the mechanical deformation varies among layers because the phases of the electric fields reflected by the load networks of adjacent PZT disks are orthogonal. Consequently, in practical implementations, the reflection signals generated within different layers exhibit a certain degree of coupling rather than being entirely independent. The influence of this coupling on the accuracy of the reflected waves will be examined through simulations and tank experiments, as detailed in Section VI.
@@ V Enhanced Matching Network
This section presents an enhanced matching network, designed in response to the varying impedance of acoustic reflectors in the dynamic underwater environment. A simplified circuit model for underwater acoustic reflectors is provided, followed by a detailed discussion on the architecture of the new matching network. We then explore various methods used to optimize its performance across different conditions.
V-A Necessity for an Enhanced Matching Network
In RF-IRS systems, the impedance of the reflector changes exclusively with frequency. Under these conditions, a wide-band matching circuit proves highly effective in aligning the impedance of the reflector with that of the load. However, for underwater acoustic reflectors constructed from PZT material, the impedance changes not only with frequency but also in response to water temperature, incident angle, and deployment depth. To ensure efficient operation of the ML-ARIS across various application scenarios, the matching circuit must be wide-band and adaptable to environmental changes.
V-A1 Impedance variation with environment and production
In Fig. 5, an Agilent E5071C vector network analyzer (VNA) is employed to investigate how the impedance of a PZT disk varies with signal frequency and water temperature. The structure of the reflector used in the test has been introduced in Fig. 3. The resonant frequencies of all disks are around 28.2 kHz at 20C. In the figure, the symbols $R$ and $X$ denote the real and imaginary components of the impedance at the first PZT disk, respectively.
@@ V-A1 Impedance variation with environment and production
As illustrated in Fig. 5, both $R$ and $X$ exhibit substantial changes at the resonant frequency with varying temperatures. For instance, at a water temperature of 9C and a signal frequency of 28.2 kHz, $R$ is 678 $\Omega$ and $X$ is 142 $\Omega$. As the temperature rises to 22C, $R$ decreases to 494 $\Omega$, and $X$ shifts to -247 $\Omega$. As a result, with the water temperature rising from 9C to 22C, the magnitude of the PZT disk impedance at 28.2 kHz increases by 20%, and the reactance of the impedance transitions from capacitive to inductive.
The impedance of the PZT material also changes with water depth. Specifically, the capacitance of PZT, which depends on their geometry and dielectric properties, is particularly sensitive to these changes. Increased water depth results in heightened pressure, which imposes mechanical stress on the material. This stress compresses the thickness of the material, consequently increasing its capacitance, which in turn leads to a reduction in the imaginary component of the impedance.
In addition to environmental effects (pressure and temperature), manufacturing and assembly tolerances will introduce unit-to-unit and layer-to-layer variability. In ML-ARIS, these production variances affect PZT disk’s dielectric/mechanical parameters and resonant frequency, as well as variations in assembly preload and bonding-layer thickness. These factors shift the effective impedance and can therefore alter the achieved reflection coefficient if the load/matching states are selected purely from nominal design values.
V-A2 Impedance variation with incident angle
For a typical RF-RIS operating in free space, the direction of incoming waves does not significantly alter the reflector’s input impedance. In contrast, for an acoustic reflector, the impedance varies noticeably with the incident angle due to changes in boundary conditions and how sound couples into or out of the piezoelectric material at different angles.
[…]
@@ V-A2 Impedance variation with incident angle
In Fig. 6, we vary the direction of incoming waves and measure the resistance and reactance for the first two layers of the PZT stack. The impedance in this figure is calculated from the ratio of the open-circuit voltage to the short-circuit current induced by the incoming waves. As illustrated, both the resistance and reactance of the PZT impedance demonstrate non-monotonic variations with changes in the incident angle. For example, at an incident angle of 90, the resistance and reactance for PZT layer 1 are 8.9 $\Omega$ and -43.5 $\Omega$, respectively. When the incident angle is reduced to 45, the resistance increases to 11.5 $\Omega$, a 29.2% increase from the initial value. Reducing the angle further to 30, the reactance decreases to -46.1 $\Omega$, a 6% reduction from the 90 incident angle.
For PZT layer 2, at an incident angle of 90, the resistance and reactance are 9.9 $\Omega$ and -42.7 $\Omega$, respectively. Changing the angle to 45 leads to an increase in resistance to 12.7 $\Omega$, marking a 28.3% rise from the 90 situation. When the angle is further decreased to 30, the reactance rises to -38.1 $\Omega$, showing a 9.7% increase compared to the 90 scenario.
@@
[…]
@@ V-C Impedance Estimation
In $\mathbf{Z_{R}}$, let $Z^{max}_{R}(f_{r})$ be the element with the maximum magnitude at the resonant frequency $f_{r}$. The parameters used in (7) to fit $Z^{max}_{R}(f)$ are represented by $R^{\alpha}_{E}$, $C^{\alpha}_{E}$, and $\left[\text{Re}^{\prime}({Z_{S}})\right]^{\alpha}$. In contrast, the impedance with the maximum magnitude at $f_{r}$ is represented by $Z^{min}_{R}(f_{r})$; the parameters used to fit $Z^{min}_{R}(f)$ are denoted as $R^{\beta}_{E}$, $C^{\beta}_{E}$, and $\left[\text{Re}^{\prime}({Z_{S}})\right]^{\beta}$, where $R^{\alpha}_{E}\!<\!R^{\beta}_{E}$, $C^{\alpha}_{E}\!<\!C^{\beta}_{E}$, and $\left[\text{Re}^{\prime}({Z_{S}})\right]^{\alpha}\!>\!\left[\text{Re}^{\prime}({Z_{S}})\right]^{\beta}$.
@@
[…]
@@ V-D1 Architecture
In Fig. 9, we illustrate the architecture of the enhanced matching circuit developed for the first layer of the PZT stack in one of the acoustic reflectors. This circuit features a three-tier subnetwork structure, each tier consisting of a high-pass L-type matching circuit. The interconnection between tiers is managed by NMOS transistors $\text{SW}_{1}$ and $\text{SW}_{2}$, which are controlled by the MCU. It is crucial to note that, the operation of the tiers is not independent but cascaded; for instance, activating tier $i$ requires that tiers 1 through i1i\!-\!1 also be activated to achieve the optimal performance.
@@ V-D2 Optimization
[…]
@@
The enhanced matching network, together with the associated optimization method, provides a mechanism for calibration-based compensation to reduce the impact of environmental variation and production-induced variability. Specifically, each PZT unit can undergo a one-time production or installation calibration using a VNA or an impedance analyzer to obtain its impedance map across the intended frequency band and expected environmental range. The resulting lookup table (or fitted model) is stored for subsequent control. During operation, the MCU selects the appropriate tier configuration and load state based on the calibrated map combined with real-time environmental estimates, thereby compensating for both environmental variability and production-induced discrepancies.
@@ VI Performance Evaluation
This section validates the design of ML-ARIS through both simulation and experimentation. We assess the performance of the enhanced matching circuit and the synthetic reflection capability of the multilayer structure. Results from tank experiments and COMSOL simulations confirm the operational viability of ML-ARIS, demonstrating its effectiveness in real-world applications.
VI-A Performance of the Enhanced Matching network
[…]
where $Z^{i}_{m}(f)$ represents the impedance of the matching circuit under environmental condition ii.
Figure 10: Comparison of the enhanced matching network and the L-type matching circuit.
As shown in Fig. 10, within the frequency band of 27.5 kHz to 28.5 kHz, the average $|\absolutevalue{\text{S}_{\text{11}}}$ using the enhanced matching network are consistently lower than those using the L-type circuit across all environmental conditions. While the L-type circuit minimizes $|\absolutevalue{\text{S}_{\text{11}}}$ at the resonant frequency, it struggles to adapt to variations in the reflector’s impedance, resulting in reduced performance in dynamic environments. In contrast, the enhanced matching network employs a cascaded structure that effectively maintains low $\text{S}_{\text{11}}$ magnitudes despite changes in the reflector’s impedance. Specifically, under environmental conditions 1, 4, and 8, the average $|\absolutevalue{\text{S}_{\text{11}}}$ values with the enhanced matching network are -21.8 dB, -15.7 dB, and -14.9 dB, respectively. These values are 3.1 dB, 2.2 dB, and 0.5 dB lower than those achieved with the L-type circuit.
Now, we connect three L-type circuits in parallel to evaluate their performance relative to the enhanced matching network. Each L-type circuit is specifically designed to match impedance within a designated range in this configuration. The final output is dynamically switched among the three circuits based on the reflector’s impedance. This contrasts with the enhanced matching network, where three-tier L-type circuits are cascaded, and the output of each tier is dependent on the input from the previous one. In the parallel matching network, the circuits function independently. The design of the parallel matching network is structured around the following optimization problem:
[…]
In Fig. 11, the resistance and capacitance of the acoustic reflector in various environments range between [200, 450] $\Omega$ and [20, 50] nF, respectively. The figure illustrates that the performance of the enhanced matching circuit with a cascaded structure is consistently better than that of the parallel configuration under all environmental conditions. Specifically, measurement data indicate that the average $|\absolutevalue{\text{S}_{\text{11}}}$ values for the enhanced and parallel matching networks are -24.1 dB and -20.7 dB, respectively, with the latter being 3.4 dB higher than the former.
Furthermore, a comparison between Fig. 11 and Fig. 10 shows that both the parallel and cascaded matching networks achieve significantly lower $|\absolutevalue{\text{S}_{\text{11}}}$ values compared to a single L-type circuit. Given that both matching networks require the same number of components (inductors, capacitors, and NMOS switches), yet the cascaded version more effectively matches the reflector’s impedance to the target load in dynamic environments, emerging as the preferred option for constructing our UA-RIS.
VI-B Verification of Multilayered Structure
VI-B1 Experimental settings
In Fig. 12, tank experiments are performed to validate the feasibility of synthetic reflection using ML-ARIS, i.e., to demonstrate that the amplitude and phase of the reflector-generated component can be regulated by adjusting the load states of individual layers. In the tests, two BTech Acoustic BT-2RCL [2] omnidirectional acoustic transducers serve as the transmitter and receiver, respectively. The transmitter is driven by a Siglent 1062X arbitrary waveform generator emitting a burst sinusoidal signal consisting of 2,500 cycles. The frequency and peak-to-peak voltage of the source signal are 43.8 kHz and 20 V, respectively. The receiver is connected to a PicoScope 4224A oscilloscope for data acquisition.